Two pipes together can fill a large tank in 10 hours. One of the pipes used alone takes 15 hours longer than the other to fill the tank. How long would each pipe take to fill the tank alone?
Answer by stanbon(57972)

Two pipes together can fill a large tank in 10 hours. One of the pipes used alone takes 15 hours longer than the other to fill the tank. How long would each pipe take to fill the tank alone?
----------------------
Together DATA:
Time = 10 hr/job : rate = 1/10 job/hr
-----------------
One tank DATA;
Time = 15 hr/job ; rate = 1/15 job/hr
--------------
Other tank DATA;
time = x hr/job ; rate = 1/x job/hr
-------------------
EQUATION:
rate + rate = together rate
1/x + 1/15 = 1/10
Multiply thru by 30x to get:
30 + 2x = 3x
x = 30 hrs (time it would take the 2nd pipe to do the job)
===========
Cheers,
Stan H.